# Fine Tuning vs. Retrieval Augmented Generation for Less Popular Knowledge

### 03 Mar 2024 v1

@@ Abstract
Large language models (LLMs) memorize a vast amount of factual knowledge, exhibiting strong performance across diverse tasks and domains. However, it has been observed that the performance diminishes when dealing with less-popular or low-frequency concepts and entities, for example in domain specific applications. The two prominent approaches to enhance the performance of LLMs on low-frequent topics are: Retrieval Augmented Generation (RAG) and fine-tuning (FT) over synthetic data. This paper explores and evaluates the impact of RAG and FT on customizing LLMs in handling low-frequency entities on question answering task. Our findings indicate that FT significantly boosts the performance across entities of varying popularity, especially in the most and least popular groups, while RAG surpasses other methods. Additionally, the success of both RAG and FT approaches is amplified by advancements in retrieval and data augmentation techniques. We release our data and code at https://github.com/HeydarSoudani/RAGvsFT.
@@
Heydar Soudani Radboud University Nijmegen The Netherlands heydar.soudani@ru.nl                        Evangelos Kanoulas University of Amsterdam Amsterdam The Netherlands e.kanoulas@uva.nl                        Faegheh Hasibi Radboud University Nijmegen The Netherlands f.hasibi@cs.ru.nl
@@ 3.2 Fine-Tuning Approach
We generate training data for the fine tuning approach using two distinct data augmentation methods. The first one is the End-to-End approach (Ushio et al., 2023), utilizing a model specifically trained for paragraph-level QA generation, using the T5 (Raffel et al., 2020) model, referred to as E2E in our paper. Additionally, we explore the generation of synthetic training data by prompting an LLM, utilizing Zephyr (Tunstall et al., 2023) for QA generation. This method is referred to as the Prompt method in our discussion. It should be noted that we generate QAs exclusively using the summary section of Wikipedia pages. This approach ensures a fair comparison between FT and RAG with ideal retrieval. Further details on the QA generation process, including the prompt, input and output examples, and statistics of the generated QAs, are provided in the Appendix B.
@@ 4 Experiments and Results
Table 2: Accuracy of base and fine-tuned models, both with and without RAG, is reported across various conditions, encompassing model size, fine-tuning approaches, and QA generation techniques. The RAG results presented are based on ideal retrieval.
Figure 3: Recall@1 performance of four retrieval methods across different popularity levels of pageviews. The results indicate that retrieval models perform more effectively with less popular knowledge compared to more popular ones.

### 07 Mar 2024 v2

@@ Abstract
Large language models (LLMs) memorize a vast amount of factual knowledge, exhibiting strong performance across diverse tasks and domains. However, it has been observed that the performance diminishes when dealing with less-popular or low-frequency concepts and entities, for example in domain specific applications. The two prominent approaches to enhance the performance of LLMs on low-frequent topics are: Retrieval Augmented Generation (RAG) and fine-tuning (FT) over synthetic data. This paper explores and evaluates the impact of RAG and FT on customizing LLMs in handling low-frequency entities on question answering task. Our findings indicate that FT significantly boosts the performance across entities of varying popularity, especially in the most and least popular groups, while RAG surpasses other methods. Additionally, the success of both RAG and FT approaches is amplified by advancements in retrieval and data augmentation techniques. The code and data is available at https://github.com/informagi/RAGvsFT.
@@
Heydar Soudani Radboud University Nijmegen The Netherlands heydar.soudani@ru.nl                        Evangelos Kanoulas University of Amsterdam Amsterdam The Netherlands e.kanoulas@uva.nl                        Faegheh Hasibi Radboud University Nijmegen The Netherlands faegheh.hasibi@ru.nl
@@ 3.2 Fine-Tuning Approach
We generate training data for the fine tuning approach using two distinct data augmentation methods. The first one is the End-to-End approach (Ushio et al., 2023), utilizing a model specifically trained for paragraph-level QA generation, using the T5 model (Raffel et al., 2020), referred to as E2E in our paper. Additionally, we explore the generation of synthetic training data by prompting an LLM, utilizing Zephyr (Tunstall et al., 2023) for QA generation. This method is referred to as the Prompt method in our discussion. We generate QAs exclusively using the summary section of Wikipedia pages to ensures a fair comparison between FT and RAG with ideal retriever. Further details on the QA generation process, including the prompt, input and output examples, and statistics of the generated QAs, are provided in the Appendix B.
@@ 4 Experiments and Results
Table 2: Accuracy of base and fine-tuned models, both with and without RAG. The RAG results presented are based on ideal retrieval.
Figure 3: Recall@1 for four retrieval models across different popularity levels. The results indicate that retrieval models perform more effectively with less popular knowledge compared to more popular ones.
@@ 6 Limitations
Acknowledgements
This publication is part of the project LESSEN with project number NWA.1389.20.183 of the research program NWA ORC 2020/21 which is (partly) financed by the Dutch Research Council (NWO).